# Robust Hallucination Detection in LLMs via Adaptive Token Selection

### v2

@@ 1 Introduction
Recent progress in Large Language Models (LLMs) has demonstrated impressive capabilities across a wide range of applications. However, the ever-growing popularity of LLMs also gives rise to concerns about the reliability of their outputs ji2023survey ; chuang2023dola . Some research has indicated that LLMs are susceptible to hallucinations, which can be described as unfaithful or incorrect generations longpre2021entity ; adlakha2023evaluating ; huang2025survey . This tendency not only impedes the broader applications of LLMs but also poses potential safety risks, especially in high-stake fields such as legal and medical services. Therefore, the reliable detection of hallucinations is critical for the safe deployment of LLMs.
@@
Various approaches have been developed to detect hallucinations. Recent studies indicate that predictive uncertainty can serve as useful detection features burns2022discovering ; kadavath2022language ; duan-etal-2024-shifting , as predictions with low confidence often correlate with the presence of hallucinated content. Some research focuses on the evaluation of a single generation manakul2023selfcheckgpt ; bakman-etal-2024-mars while some harness the information contained in multiple generations. Compared to single-generation methods, approaches utilising multiple generations demonstrate greater effectiveness since they capture a broader set of hallucination signals. Such signals include semantic consistency farquhar2024detecting for the repeated queries, as well as behaviour inconsistency when LLMs are presented with a set of unrelated follow-up questions pacchiardi2024how after the initial question. Nevertheless, these methods are mainly based on the final generation output, rendering them ineffective in leveraging important semantics in the internal representations.
@@
In parallel, another line of research focuses on the utilisation of the internal representations of LLMs for hallucination detection azaria-mitchell-2023-internal ; liu2025attention . These internal representations can encode information about the truthfulness direction of the generations. Motivated by this, great efforts have been made to explore the characteristics of these representations, most commonly by training a binary classifier on them. Various supervision signals, including accuracy labels li2024inference , converted semantic entropy labels kossen2024semantic , and eigenvalue-related labels du2024haloscope , have been harnessed to train the classifier for detection tasks. One major challenge for these methods is that a majority of tokens in an incorrect/hallucinated response may not contribute to truthfulness. To address this issue, most methods utilise predetermined tokens, such as the first generated token, the last generated token, or the one before the last ji-etal-2024-llm ; barbero2025llms ; zucchet2025language . However, the exact location of the most indicative tokens for hallucination can vary significantly, as illustrated in Figure 1, since the generation responses are of free-form with varying lengths and have sparse distributions of hallucinated entities. As a result, they can overlook important tokens where hallucinated information is actually concentrated.
@@
To address this challenge, we propose HaMI, namely, Hallucination detection as Multiple Instance learning, a novel approach that jointly optimises token selection and hallucination detection in an end-to-end fashion. The joint optimisation enables adaptive token selection from internal state representations for stable and accurate hallucination detection on generation responses of varying length. In HaMI we reformulate the task as a multiple instance learning (MIL) problem carbonneau2018multiple , where each response sequence is treated as a bag of token instances, with a bag-level label as either hallucinated (positive) or trustworthy (negative), and the objective becomes binary classification of the token bags. This way takes advantage of the fact that only a few token instances in the positive bag are positive since hallucinated content typically manifests in only a small subset of tokens within a sequence, whereas all token instances in the negative bag are always negative. In doing so, the MIL approach enables the exploitation of the hallucination labels at the sequence (bag) level to adaptively select the most responsive tokens for sequence-level hallucination detection.
To be more specific, LLMs are first prompted to generate response sequences with varying length. The MIL-driven hallucination detector is then optimised in HaMI to assign hallucination scores to all individual token instances and adaptively select the most indicative tokens in both positive and negative bags for the sequence-level prediction. The optimisation results in a detector that can distinguish the most positive hallucinated token instances from the hard negative token instances (i.e.,, the tokens that have highest hallucination scores in a negative bag). Additionally, recognizing that predictive uncertainty serves as an important indicator of correctness, we further propose a representation enhancement module in HaMI, where we integrate multiple levels of uncertainty information into the original representation space for more effective training of our HaMI detector.
@@ 2 Related Work
The term hallucination, under the closed-book setting, can refer to unfaithful or fabricated generations ji2023survey ; zhang2023siren . Having gained wide interest, hallucination detection is crucial for LLMs to maintain high reliability in various specific tasks. These methods can be categorised into two main lines: uncertainty measurement and internal representation analysis.
@@
Uncertainty measurement has been widely explored for hallucination detection. Some research focuses on token-level uncertainty talman-etal-2023-uncertainty ; duan-etal-2024-shifting ; quevedo2024detecting with the assumption that low predictive logits or high entropy over tokens’ predictive distribution indicate a high possibility of hallucination. Some studies target sentence-level uncertainty to capture a more holistic view of the whole response. Such uncertainty can be measured by (weighted) aggregation of all token logits bakman-etal-2024-mars ; ren2023outofdistribution or obtained by directly instructing LLMs to express the truthfulness of their generations with simple but effective prompts kadavath2022language ; lin2022teaching ; zhou2023navigating ; manakul2023selfcheckgpt . To further enrich semantic understanding and improve the detection accuracy, researchers also explore LLMs’ behaviour among multiple generations mundler2023self ; dhuliawala2023chain . For example, Farquhar et al. farquhar2024detecting propose Semantic Entropy as an evaluation of semantic consistency among multiple generations for hallucination detection, while Pacchiardi et al. pacchiardi2024how present that behaviour inconsistency over predefined follow-up unrelated questions can also indicate the correctness of the response to the initial question. Although effective, the performance of these methods relies on external tools and ignores the important semantics embedded in the internal representations of LLMs.
@@
Recently, a branch of work suggests that the internal representations of LLMs encode more knowledge than they express and can reveal truthfulness direction yuksekgonul2023attention ; hubinger2024sleeper ; snyder2024early ; bayat2025steering . The majority of this line employs probes alain2016understanding to better understand head-wise or layer-wise representations and predict the correctness of generations burns2022discovering ; li2024inference ; chen2024inside ; marks2024geometrytruthemergentlinear . Recent research extends these methods by proposing new supervision signals, such as an automated membership estimation score presented in HaloScope du2024haloscope and the aforementioned semantic entropy value kossen2024semantic , which has been proven to be preferable to accuracy labels for supervised training. Most of these works leverage predefined token representations, but the truthfulness information is concentrated in specific tokens. Some research attempts to prompt an LLM to find the “exact token” in the sequence orgad2024llmsknowshowintrinsic , but the reliability of detection is greatly influenced by the capability of the employed LLM. Unlike existing methods that separate token selection and hallucination detection into two stages and require external assistance, we propose to train a ranking model to automatically select the critical tokens for hallucination detection in an end-to-end manner.
@@ 3 Preliminaries
Given a sequence of input tokens $\mathbf{x}=\{x_{1},x_{2},\dots,x_{m}\}$ consisting of a specific question $q$, an LLM will generate a sequence of tokens $\mathbf{y}=\{y_{1},y_{2},\dots,y_{t}\}$. Generally, each token $y_{i\in\{1,2,\dots,t\}}$ is decoded from the next predictive distribution over the model’s vocabulary set $\mathcal{V}$, formulated as $y_{i}=\arg\max_{y\in\mathcal{V}}P(y|y_{<i},\mathbf{x})$, and the predictive probability is denoted by $p_{i}$ for short. By accessing the internal states of the model, we can extract the internal representation $h_{i,l}\in\mathbbm{R}^{d}$ at layer $l$ for each token $y_{i}$, where $d$ is decided by the dimensions of the internal states of the LLM. The correctness of generated response is evaluated by GPT-4.1 achiam2023gpt with label $z\in\{0,1\}$.
Given a datset $D=\{(q_{n},a_{n})\}_{n=1}^{N}$, where $\{q_{n}\}_{n=1}^{N}$ and $\{a_{n}\}_{n=1}^{N}$ are questions and ground-truth answers respectively, LLMs will generate answers decoded from a token list $\mathbf{y}_{n}$ accompanied with predictive probabilities, hidden representations $\mathbf{h}_{n}$ and correctness labels $z_{n}\in\{0,1\}$. The representation space for all samples is denoted as =$\mathcal{H}=\{\mathbf{h}_{n}\}_{n=1}^{N}$ Our MIL-based hallucination detection is to identify the most representative tokens in incorrect responses and the hard negative tokens (i.e., the most likely hallucinated tokens within correct responses) that are optimal for training the subsequent detector, i.e.,, a joint optimisation of token selection and hallucination detection.
@@ 4.1 MIL-driven Adaptive Token Selection
For a given sentence, its correctness is often determined by just a few words, such as noun entities chuang2023dola , suggesting that truthfulness information may be encoded in the internal representations of specific tokens. Nevertheless, the correctness label is typically assigned to the entire sequence. The key idea of our approach is to adaptively identify the most salient tokens, upon which we can train a reliable hallucination detector. To this end, we introduce Multiple Instance Learning (MIL) to this domain.
@@
Therefore, we reformulate hallucination detection with adaptive token selection as a MIL problem. Particularly, the generated sequence can be regarded as a bag of tokens. Generations without hallucinations are labelled as negative bags $\mathcal{B^{-}}$ (label ‘0’), while those with hallucinations are labelled as positive bags $\mathcal{B^{+}}$ (label ‘1’). The representations of the positive and negative sequences are denoted as $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$ respectively and $\mathbf{h}=\{h_{i}\}_{i=1}^{t}$. Intuitively, the detector is expected to assign higher scores to token instances from the positive bag compared to those from the negative bag. Given the hallucination sparsity insight mentioned before, we choose the token instances with the top $k$ highest hallucination scores as the salient tokens in each sequence, which can be defined as
@@
where $i_{k}\in\{1,2,\cdots,t\}$, and $f_{\theta}$ is the hallucination detector we aim to learn with the parameters $\theta$, as presented in Figure 2. As such, we are able to locate the most hallucinated token instances in the positive bag and the top hard negative token instances resembling a hallucination in the negative bag as the salient tokens. Our approach then seeks to distinguish between positive and negative bags by maximizing the distance between the selected token instances from these two categories in the representation space. The training MIL objective can be formulated as follows:
@@
where $i\in\{2,\cdots,t\}$. Through this, we aim to ensure the consistency of the hallucination scores of neighbouring tokens. Therefore, HaMI is optimised by minimizing the following overall loss:
@@ 4.2 Augmenting Internal State Representations with Predictive Uncertainty
where $\mathbf{x}$ are prompt tokens, $y$ are generated tokens and $i\in\{1,2,\cdots,t\}$; ii) sentence-level perplexity, which is monotonically related to the mean of the negative log-likelihood of the sequence:
@@
and iii) semantic consistency across multiple samples, where the uncertainty value can be quantified by the number of semantic-equivalence generations over the whole generations based on the entailment results. We choose Semantic Entropy here farquhar2024detecting . Specifically,
@@ 5 Experiments
We evaluate our method on four popular benchmark datasets across a range of question-answering (QA) domains, including (1) Trivia QA joshi2017triviaqa , a relatively complicated confabulation QA datasets; (2) Stanford Question Answering Dataset (SQuAD for short) rajpurkar2016squad , which is based on Wikipedia and generated by humans through crowdsourcing; (3) Natural Questions (denoted as NQ) kwiatkowski2019natural , containing search information from real users on Google search and (4) a biomedical QA corpus BioASQ krithara2023bioasq . For each dataset, we randomly extract $2,000$ QA pairs for training and $800$ pairs for testing. For multi-sampling approaches, we prompt LLMs six times for each question for the test set generation. Following farquhar2024detecting , we have a total of $500$ questions randomly sampled from the generated test set, with approximately $400$ questions retained as the final inputs at the testing stage for evaluation. The refined set from the remaining $300$ QA pairs is used as a validation set for selecting the optimal layer for representation extraction. We employ representative open-sourced LLMs, LLaMA touvron2023llama ; dubey2024llama and Mistral family jiang2023mistral , and evaluate our approach on LLaMA-3.1-8B and Mistral-Nemo-Instruct (12B), as well as a larger version of LLaMA, namely LLaMA-3.13-Instruct-70B, which is deployed with 8-bit quantisation wolf-etal-2020-transformers . All answers are generated under the temperature of $0.5$ and with context-free zero-shot prompts. Please see more details for the setup in Appendix.
@@
We compare our method with a series of state-of-the-art (SotA) methods, covering both uncertainty-based methods and internal representation-based methods. The uncertainty-based methods include: p(True) kadavath2022language , that asks LLMs to express the correctness for given answers themselves; Perplexity ren2023outofdistribution , which is based on Eq. 8; Semantic Entropy (SE) farquhar2024detecting , semantic-equivalence measurement across multiple samples, where GPT-3.5 is used for entailment evaluation; MARS bakman-etal-2024-mars , that performs a weighted aggregation of token logits, and its enhanced version (MARS-SE), augmented with semantic entropy. For internal representation-based methods, probing classifiers with various supervision signals are included: CCS burns2022discovering , utilising contrast-consistent discovering for correctness probability measurements; SAPLMA azaria-mitchell-2023-internal , trained with correctness labels; HaloScope du2024haloscope , which proposes an automated membership estimation score and converts the score to binary labels for classification. All these methods are based on the last generated token as settled in their original paper. Notably, SAPLMA utilises a four-layer multilayer perception (MLP) with ReLU non-linearity while the other methods, including HaMI, employ a detector with two layers. We also assess the performance of HaMI using varying number of layers in Section 5.3.
@@
Following farquhar2024detecting ; ren2023outofdistribution , we evaluate the model’s capability for hallucination detection by calculating the area under the receiver operating characteristic curve (AUROC). The ground-truth labels indicating the correctness are given by GPT-4.1 achiam2023gpt , which is instructed to determine if the answer is correct or not based on the consistency between generated answers and gold answers and their own knowledge. The prompt template follows farquhar2024detecting (see more details in Appendix). Note that since GPT-4.1 makes mistakes mostly for the positive samples, we ask GPT-4.1 to rejudge samples labelled as positive and discard samples if the result is inconsistent with the first result. For this reason, the final sizes of the training and testing sets vary around $1,900$ and $400$, respectively.

### v3

@@ 1 Introduction
Recent progress in Large Language Models (LLMs) has demonstrated impressive capabilities across a wide range of applications. However, the ever-growing popularity of LLMs also gives rise to concerns about the reliability of their outputs [22, 12]. Some research has indicated that LLMs are susceptible to hallucinations, which can be described as unfaithful or incorrect generations [33, 2, 18]. This tendency not only impedes the broader applications of LLMs but also poses potential safety risks, especially in high-stake fields such as legal and medical services. Therefore, the reliable detection of hallucinations is critical for the safe deployment of LLMs.
@@
Various approaches have been developed to detect hallucinations. Recent studies indicate that predictive uncertainty can serve as useful detection features [9, 25, 15], as predictions with low confidence often correlate with the presence of hallucinated content. Some research focuses on the evaluation of a single generation [34, 6] while some harness the information contained in multiple generations. Compared to single-generation methods, approaches utilising multiple generations demonstrate greater effectiveness since they capture a broader set of hallucination signals. Such signals include semantic consistency [17] for the repeated queries, as well as behaviour inconsistency when LLMs are presented with a set of unrelated follow-up questions [38] after the initial question. Nevertheless, these methods are mainly based on the final generation output, rendering them ineffective in leveraging important semantics in the internal representations.
@@
In parallel, another line of research focuses on the utilisation of the internal representations of LLMs for hallucination detection [5, 32]. These internal representations can encode information about the truthfulness direction of the generations. Motivated by this, great efforts have been made to explore the characteristics of these representations, most commonly by training a binary classifier on them. Various supervision signals, including accuracy labels [30], converted semantic entropy labels [26], and eigenvalue-related labels [14], have been harnessed to train the classifier for detection tasks. One major challenge for these methods is that a majority of tokens in an incorrect/hallucinated response may not contribute to truthfulness. To address this issue, most methods utilise predetermined tokens, such as the first generated token, the last generated token, or the one before the last [21, 7, 52]. However, the exact location of the most indicative tokens for hallucination can vary significantly, as illustrated in Figure 1, since the generation responses are of free-form with varying lengths and have sparse distributions of hallucinated entities. As a result, they can overlook important tokens where hallucinated information is actually concentrated.
@@
To address this challenge, we propose HaMI, namely, Hallucination detection as Multiple Instance learning, a novel approach that jointly optimises token selection and hallucination detection in an end-to-end fashion. The joint optimisation enables adaptive token selection from internal state representations for stable and accurate hallucination detection on generation responses of varying length. In HaMI, we reformulate the task as a multiple instance learning (MIL) problem [10], where each response sequence is treated as a bag of token instances, with a bag-level label as either hallucinated (positive) or trustworthy (negative), and the objective becomes binary classification of the token bags. This way takes advantage of the fact that only a few token instances in the positive bag are positive since hallucinated content typically manifests in only a small subset of tokens within a sequence, whereas all token instances in the negative bag are always negative. In doing so, the MIL approach enables the exploitation of the hallucination labels at the sequence (bag) level to adaptively select the most responsive tokens for sequence-level hallucination detection.
To be more specific, LLMs are first prompted to generate response sequences with varying length. The MIL-driven hallucination detector is then optimised in HaMI to assign hallucination scores to all individual token instances and adaptively select the most indicative tokens in both positive and negative bags for the sequence-level prediction. The optimisation results in a detector that can distinguish the most positive hallucinated token instances from the hard negative token instances (i.e., the tokens that have the highest hallucination scores in a negative bag). Additionally, recognising that predictive uncertainty serves as an important indicator of correctness, we further propose a representation enhancement module in HaMI, where we integrate multiple levels of uncertainty information into the original representation space for more effective training of our HaMI detector.
@@ 2 Related Work
The term hallucination, under the closed-book setting, can refer to unfaithful or fabricated generations [22, 50]. Having gained wide interest, hallucination detection is crucial for LLMs to maintain high reliability in various specific tasks. These methods can be categorised into two main lines: uncertainty measurement and internal representation analysis.
@@
Uncertainty measurement has been widely explored for hallucination detection. Some research focuses on token-level uncertainty [46, 15, 40] with the assumption that low predictive logits or high entropy over tokens’ predictive distribution indicate a high possibility of hallucination. Some studies target sentence-level uncertainty to capture a more holistic view of the whole response. The uncertainty can be measured by (weighted) aggregation of all token logits, such as Perplexity [42], Mars [6] and G-NLL [3], or obtained by directly instructing LLMs to express the truthfulness of their generations with simple but effective prompts [25, 31, 51, 34].
To further enrich semantic understanding and improve the detection accuracy, researchers also explore LLMs’ behaviour among multiple generations [36, 13]. For example, Farquhar et al. [17] propose Semantic Entropy as an evaluation of semantic consistency among multiple generations for hallucination detection, while Pacchiardi et al. [38] present that behaviour inconsistency over predefined follow-up unrelated questions can also indicate the correctness of the response to the initial question. Although effective, the performance of these methods relies on external tools and ignores the important semantics embedded in the internal representations of LLMs.
@@
Recently, a branch of work suggests that the internal representations of LLMs encode more knowledge than they express and can reveal truthfulness direction [49, 19, 44, 8]. The majority of this line employs probes [4] to better understand head-wise or layer-wise representations and predict the correctness of generations [9, 30, 11, 35]. Recent research extends these methods by proposing new supervision signals, such as an automated membership estimation score presented in HaloScope [14] and the aforementioned semantic entropy value [26], which has been proven to be preferable to accuracy labels for supervised training. Most of these works leverage predefined token representations, but the truthfulness information is concentrated in specific tokens. Some research attempts to prompt an LLM to find the “exact token” in the sequence [37], but the reliability of detection is greatly influenced by the capability of the employed LLM. Unlike existing methods that separate token selection and hallucination detection into two stages and require external assistance, we propose to train a ranking model to automatically select the critical tokens for hallucination detection in an end-to-end manner.
@@ 3 Preliminaries
Given a sequence of input tokens $\mathbf{x}=\{x_{1},x_{2},\dots,x_{m}\}$ consisting of a specific question $q$, an LLM will generate a sequence of tokens $\mathbf{y}=\{y_{1},y_{2},\dots,y_{t}\}$. Generally, each token $y_{i\in\{1,2,\dots,t\}}$ is decoded from the next predictive distribution over the model’s vocabulary set $\mathcal{V}$, formulated as $y_{i}=\arg\max_{y\in\mathcal{V}}P(y|y_{<i},\mathbf{x})$, and the predictive probability is denoted by $p_{i}$ for short. By accessing the internal states of the model, we can extract the internal representation $h_{i,l}\in\mathbbm{R}^{d}$ at layer $l$ for each token $y_{i}$, where $d$ is decided by the dimensions of the internal states of the LLM. The correctness of generated response is evaluated by GPT-4.1 [1] with label $z\in\{0,1\}$.
Given a datset $D=\{(q_{n},a_{n})\}_{n=1}^{N}$, where $\{q_{n}\}_{n=1}^{N}$ and $\{a_{n}\}_{n=1}^{N}$ are questions and ground-truth answers respectively, LLMs will generate answers decoded from a token list $\mathbf{y}_{n}$ accompanied with predictive probabilities, hidden representations $\mathbf{h}_{n}$ and correctness labels $z_{n}\in\{0,1\}$. The representation space for all samples is denoted as $\mathcal{H}=\{\mathbf{h}_{n}\}_{n=1}^{N}$ Our MIL-based hallucination detection is to identify the most representative tokens in incorrect responses and the hard negative tokens (i.e., the most likely hallucinated tokens within correct responses) that are optimal for training the subsequent detector, i.e., a joint optimisation of token selection and hallucination detection.
@@ 4.1 MIL-driven Adaptive Token Selection
For a given sentence, its correctness is often determined by just a few words, such as noun entities [12], suggesting that truthfulness information may be encoded in the internal representations of specific tokens. Nevertheless, the correctness label is typically assigned to the entire sequence. The key idea of our approach is to adaptively identify the most salient tokens, upon which we can train a reliable hallucination detector. To this end, we introduce Multiple Instance Learning (MIL) to this domain.
@@
Therefore, we reformulate hallucination detection with adaptive token selection as an MIL problem. Particularly, the generated sequence can be regarded as a bag of tokens. Generations without hallucinations are labelled as negative bags $\mathcal{B^{-}}$ (label ‘0’), while those with hallucinations are labelled as positive bags $\mathcal{B^{+}}$ (label ‘1’). The representations of the positive and negative sequences are denoted as $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$ respectively and $\mathbf{h}=\{h_{i}\}_{i=1}^{t}$. Intuitively, the detector is expected to assign higher scores to token instances from the positive bag compared to those from the negative bag. Given the hallucination sparsity insight mentioned before, we choose the token instances with the top $k$ highest hallucination scores as the salient tokens in each sequence, which can be defined as
@@
where $i_{k}\in\{1,2,\cdots,t\}$, and $f_{\theta}$ is the hallucination detector we aim to learn with the parameters $\theta$, as presented in Figure 2. As such, we are able to locate the most hallucinated token instances in the positive bag and the top hard negative token instances resembling a hallucination in the negative bag as the salient tokens. Our approach then seeks to distinguish between positive and negative bags by maximising the distance between the selected token instances from these two categories in the representation space. The training MIL objective can be formulated as follows:
@@
where $i\in\{2,\cdots,t\}$. Through this, we aim to ensure the consistency of the hallucination scores of neighbouring tokens. The effectiveness of $\mathcal{L}_{Smooth}$ is discussed in Appendix B in detail. Finally, HaMI is optimised by minimising the following overall loss:
@@ 4.2 Augmenting Internal State Representations with Predictive Uncertainty
where $\mathbf{x}$ are prompt tokens, $y$ are generated tokens and $\in\{1,2,\cdots,t\}; ii$) sentence-level perplexity, which is monotonically related to the mean of the negative log-likelihood of the sequence:
@@
and iii) semantic consistency across multiple samples, where the uncertainty value can be quantified by the number of semantic-equivalence generations over the whole generations based on the entailment results. We choose Semantic Entropy here [17]. Specifically,
@@ 5 Experiments
We evaluate our method on four popular benchmark datasets across a range of question-answering (QA) domains, including (1) Trivia QA [24], a relatively complicated confabulation QA datasets; (2) Stanford Question Answering Dataset (SQuAD for short) [41], which is based on Wikipedia and generated by humans through crowdsourcing; (3) Natural Questions (denoted as NQ) [28], containing search information from real users on Google search and (4) a biomedical QA corpus BioASQ [27]. For each dataset, we randomly extract $2,000$ QA pairs for training and $800$ pairs for testing. For multi-sampling approaches, we prompt LLMs six times for each question for the test set generation. Following [17], we have a total of $500$ questions randomly sampled from the generated test set, with approximately $400$ questions retained as the final inputs at the testing stage for evaluation. The refined set from the remaining $300$ QA pairs is used as a validation set for selecting the optimal layer for representation extraction. We employ representative open-sourced LLMs, LLaMA [47, 16] and Mistral family [23], and evaluate our approach on LLaMA-3.1-8B and Mistral-Nemo-Instruct (12B), as well as a larger version of LLaMA, namely LLaMA-3.3-Instruct-70B, which is deployed with 8-bit quantisation [48]. All answers are generated under the temperature of $0.5$ and with context-free zero-shot prompts. Please see more details for the setup in Appendix A.2.
@@
We compare our method with a series of state-of-the-art (SotA) methods, covering both uncertainty-based methods and internal representation-based methods. The uncertainty-based methods include: p(True) [25], that asks LLMs to express the correctness of given answers themselves; Perplexity [42], which is based on Eq. 8; Semantic Entropy (SE) [17], semantic-equivalence measurement across multiple samples, where GPT-3.5 is used for entailment evaluation; MARS [6], that performs a weighted aggregation of token logits, and its enhanced version (MARS-SE), augmented with semantic entropy. For internal representation-based methods, probing classifiers with various supervision signals are included: CCS [9], utilising contrast-consistent discovering for correctness probability measurements; SAPLMA [5], trained with correctness labels; HaloScope [14], which proposes an automated membership estimation score and converts the score to binary labels for classification; CED [29], exploiting specially designed auxiliary and oracle samples to enhance the distinction between in-distribution and out-of-distribution data; and LLM-Check [45], employing token logits and eigenvalue analysis of internal representation for detection, from which we choose the best-performing hidden state (LLM-Check-h) and attention-based variations (LLM-Check-a). All these methods are based on the last generated token as settled in their original paper. Notably, SAPLMA utilizes a four-layer multilayer perception (MLP) with ReLU non-linearity while the other methods, including HaMI, employ a detector with two layers. We also assess the performance of HaMI using varying number of layers in Section 5.3.
@@
Following [17, 42], we evaluate the model’s capability for hallucination detection by calculating the area under the receiver operating characteristic curve (AUROC). The ground-truth labels indicating the correctness are given by GPT-4.1 [1], which is instructed to determine if the answer is correct or not based on the consistency between generated answers and gold answers and its own knowledge. The prompt template follows [17] (see more details in Appendix A.2). Note that since GPT-4.1 makes mistakes mostly for the positive samples, we ask GPT-4.1 to rejudge samples labelled as positive and discard samples if the result is inconsistent with the first result. For this reason, the final sizes of the training and testing sets vary around $1,900$ and $400$, respectively. Additionally, we also conduct experiments where answers are evaluated with BLEURT score[43]. Results are presented in Appendix D.